# Kernel Ridge Regression for Efficient Learning of High-Capacity Hopfield Networks

### v1

@@
[…]
Kernel Ridge Regression for Efficient Learning of High-Capacity Hopfield Networks
(2015; 2015)
keywords:
Hopfield network, Kernel ridge regression, Associative memory, Storage capacity, Closed-form solution, Efficient learning.
{summary}
Hebbian learning limits Hopfield network capacity. While kernel methods like Kernel Logistic Regression (KLR) improve performance via iterative learning, we propose Kernel Ridge Regression (KRR) as an alternative. KRR learns dual variables non-iteratively via a closed-form solution, offering significant learning speed advantages. We show KRR achieves comparably high storage capacity (reaching ratio 1.5 shown) and noise robustness (recalling from around 80% corrupted patterns) as KLR, while drastically reducing training time, establishing KRR as an efficient method for building high-performance associative memories.
@@ 1 Introduction
Hopfield networks [1] model content-addressable memory but suffer from limited storage capacity ($\beta=P/N\approx 0.14$, where $P$ is the number of patterns and $N$ is the number of neurons) when using the standard Hebbian rule [2], leading to retrieval errors for higher loads. Various approaches have sought to improve capacity, often framing learning as a supervised task to ensure patterns are stable fixed points. While linear methods like LLR [3] offer moderate improvements, they are constrained by linear separability. Non-linear kernel methods show greater promise; Kernel Logistic Regression (KLR) recently demonstrated dramatically enhanced capacity and robustness via iterative learning. However, this iterative process can be computationally intensive [4].
In this letter, we propose Kernel Ridge Regression (KRR) [5] as an alternative, efficient kernel-based approach for Hopfield networks. KRR utilizes the kernel trick but predicts bipolar states via regression with a squared-error loss, crucially offering a non-iterative, closed-form learning solution. We implement and evaluate KRR, comparing its capacity and robustness against Hebbian, LLR, and KLR. Our aim is to verify if KRR can match KLR’s high performance while offering substantial training speed advantages. Our results confirm KRR achieves comparable state-of-the-art performance with significantly reduced learning time, establishing it as a potent and efficient training method.
This work provides the first (or among the first) empirical demonstration comparing KRR with KLR in the context of Hopfield networks, highlighting KRR’s significant efficiency gains.
@@ 2.1 Model Setup
We consider a network of $N$ bipolar neurons $\{-1,1\}^{N}$. Let $\{\boldsymbol{\xi}^{\mu}\}_{\mu=1}^{P}$ be the set of $P$ bipolar patterns to be stored in the network’s associative memory. For supervised learning approaches requiring binary targets (LLR and KLR), we transform these patterns into target vectors $\mathbf{t}^{\mu}\in\{0,1\}^{N}$ where $t_{i}^{\mu}=(\xi_{i}^{\mu}+1)/2$. For KRR, the original bipolar values $\xi_{i}^{\mu}$ are used as targets. The network state at discrete time $t$ is represented by the vector $\mathbf{s}(t)\in\{-1,1\}^{N}$. This state evolves over subsequent time steps according to the dynamics determined by the specific learning algorithm employed (detailed in Sec. 2.2 and 2.3).
@@ 2.2.1 Baselines
Hebbian Learning:
The standard Hebbian rule [1, 2] computes the weight matrix $\mathbf{W}^{\text{Heb}}=(1/N)\mathbf{X}^{\top}\mathbf{X}$ (diagonals zeroed). Recall uses $\mathrm{s}_{i}(t+1)=\text{sign}(\sum_{j\neq i}\mathbf{W}_{ij}^{\text{Heb}}\mathbf{s}_{j}(t))$ This serves as the classical baseline.
Linear Logistic Regression (LLR):
A linear supervised baseline [3], where weights $\mathbf{W}^{\text{LLR}}$ are learned iteratively via gradient descent to optimize logistic loss with a linear predictor for predicting neuron states. Recall uses the same update rule as Hebbian, with $\mathbf{W}^{\text{LLR}}$.
Kernel Logistic Regression (KLR):
An iterative kernel-based method shown to achieve high capacity [4]. It learns dual variables $\boldsymbol{\alpha}^{\text{KLR}}$ by iteratively optimizing regularized logistic loss. Recall involves kernel evaluations (Sec. 2.3). We use results from [4] for comparison.
@@ 2.3 Recall Process
Hebbian and LLR:
$s_{i}(t)=\text{sign}(\sum_{j\neq i}\mathbf{W}_{ij}s_{j}(t))$ (using $\mathbf{W}^{\text{Heb}}$ or $\mathbf{W}^{\text{Heb}}$), where $s_{i}(t)$ is the $i$-th element of the state $\mathbf{s}(t)$.
KLR and KRR:
Recall operates without an explicit $N\times N$ weight matrix $\mathbf{W}$, instead utilizing the learned $P\times N$ dual variable matrix $\boldsymbol{\alpha}$ and the stored patterns $\mathbf{X}$. The update rule involves three steps, performed synchronously for all neurons:
@@ 3.1 Experimental Setup
We conducted simulations to evaluate the performance of the proposed KRR approach and compare it against baseline methods. All simulations utilized networks composed of $N=500$ bipolar neurons. Random bipolar patterns $\boldsymbol{\xi}^{\mu}\in\{0,1\}^{N}$ were generated for storage, with each element chosen independently and with equal probability $P(\xi_{i}^{\mu}=1)=0.5$.
We compared four learning algorithms:
Hebbian:
Standard Hebbian learning rule [1, 2].
LLR:
Linear Logistic Regression, implemented as described in [3], serving as a linear supervised baseline.
KLR:
Kernel Logistic Regression, using the implementation and results reported in our previous work [4] for direct comparison.
KRR:
The proposed Kernel Ridge Regression approach (Sec. 2.2.2).
For both KLR and KRR, we employed the Radial Basis Function (RBF) kernel $K(\mathbf{x},\mathbf{y})=\exp(\gamma\|\mathbf{x}-\mathbf{y}\|^{2})$ with the kernel width parameter set to $\gamma=1/N$. Based on preliminary experiments and findings in [4], the L2 regularization parameter was set to $\lambda=0.01$ for both LLR, KLR and KRR, as this value provided robust performance. For the iterative learning methods (LLR and KLR), the learning rate was $\eta=0.1$ and the number of learning updates was fixed at 200, ensuring sufficient convergence in our previous study [4].
@@ 3.2 Storage Capacity and Noise Robustness Evaluation
We first assessed the storage capacity of each learning algorithm by measuring the recall success rate as a function of the storage load $\beta=P/N$. Recall was initiated from the original stored patterns ($\mathbf{s}(0)=\boldsymbol{\xi}^{\mu}$), testing the network’s ability to maintain them as stable fixed points. Figure 1 plots the success rate against $\beta$ for $N=500$ neurons. As expected, the standard Hebbian network exhibits a sharp performance drop around the theoretical limit of $\beta\approx 0.14$ [2]. LLR significantly improves upon this, maintaining near-perfect recall up to $\beta\approx 0.85$, but fails completely as the load approaches $\beta=0.95$. In stark contrast, both KLR, consistent with our previous findings [4], and the proposed KRR demonstrate substantially higher capacity. Both kernel methods achieved a 100% success rate across the entire tested range, including loads exceeding $\beta=1.0$ (up to $\beta=1.5$ shown), indicating their capability to reliably store patterns even when $P>N$.
Next, we evaluated noise robustness by examining the size of the basins of attraction. We fixed the storage load at an intermediate level, $\beta=0.2$ ($P=100$ for $N=500$), where both LLR and kernel methods typically succeed from clean initial states, but Hebbian learning fails. We initiated recall from corrupted versions of a target pattern, varying the initial overlap $m(0)$. Figure 2 shows the mean final overlap $m(T)$ achieved after $T=25$ steps as a function of $m(0)$. The Hebbian network fails to converge to the target pattern regardless of the initial overlap within this range. LLR shows improved robustness, successfully recalling the pattern ($m(T)\approx 1.0$) if the initial overlap $m(0)$ is above approximately 0.4. Both KLR and KRR exhibit significantly enhanced robustness. They achieve perfect recall ($m(T)\approx 1.0$) even when starting from much noisier states, requiring an initial overlap $m(0)$ of only about 0.2. This demonstrates that both kernel methods create considerably larger basins of attraction compared to LLR and, especially, the Hebbian rule, making them more resilient to input noise. Notably, the robustness profiles of KLR and KRR appear virtually identical under these conditions.
@@
Figure 1: Recall success rate vs. storage load ($\beta=P/N$) for Hebbian, LLR, and KRR ($N=500$).
@@
Figure 2: Noise robustness comparison: Final Overlap $m(T)$ vs. Initial Overlap $m(0)$ at $\beta=0.2$ for Hebbian, LLR, and KRR.
@@ 3.3 Learning Time Comparison
To quantitatively assess the practical efficiency trade-offs alongside the performance benefits, we measured and compared the learning phase computational time for LLR, KLR, and KRR across various storage loads ($\beta$). Wall-clock time was measured using Python’s time.time() function, capturing the time required to complete the learning process for each method on a standard desktop CPU. The results, averaged over 3 trials for $N=500$, are presented in Table 1 and visualized in Figure 3 (note the logarithmic scale for the time axis).
@@
Hebbian learning, being non-iterative and computationally simple ($O(N^{2})$), completes extremely quickly. Its learning time is omitted from the comparison table/figure for clarity, as it is significantly shorter than the other methods. The timing results reveal significant differences in efficiency and scaling among the supervised methods. LLR, requiring 100 iterative updates in our setup, exhibited the slowest performance, with learning time increasing approximately linearly with the storage load $\beta$ (from around 3 seconds at $\beta=0.1$ to around 30 seconds at $\beta=1.0$. This scaling is consistent with its need to process $P$ patterns across $N$ neurons iteratively.
@@
KLR, using 200 iterative updates, was considerably faster than LLR across all tested loads. Its learning time, however, still showed a clear dependence on $\beta$, increasing from around 0.07 seconds at $\beta=0.1$ to around 1.35 seconds at $\beta=1.0$. This increase is likely dominated by computations involving the $P\times P$ kernel matrix within the iterative optimization process.
@@
Crucially, KRR demonstrated outstanding learning efficiency. Leveraging its non-iterative, closed-form solution (Eq. 1), KRR consistently achieved the fastest learning times by a large margin. Its learning time remained under 0.1 seconds even at the maximum load of $\beta=1.0$ ( 0.07 seconds), showing only a very modest increase with $\beta$. This time is primarily composed of the one-time kernel matrix construction and the subsequent solving of the linear system. At $\beta=1.0$, KRR was approximately 20 times faster than KLR and over 400 times faster than LLR, highlighting the dramatic speed advantage conferred by its closed-form solution.
These results underscore the significant practical benefit of KRR for training Hopfield networks. While both KLR and KRR deliver comparable, state-of-the-art storage performance, KRR offers substantially higher training efficiency, especially at higher storage loads. This makes KRR a highly attractive option, particularly when computational resources for training are limited or when models need to be trained rapidly. The scalability challenge for recall ($O(PN)$ complexity), however, remains similar for both KLR and KRR.

### v4

@@
Abstract
Hopfield networks using Hebbian learning suffer from limited storage capacity. While supervised methods like Linear Logistic Regression (LLR) offer some improvement, kernel methods like Kernel Logistic Regression (KLR) significantly enhance storage capacity and noise robustness. However, KLR requires computationally expensive iterative learning. We propose Kernel Ridge Regression (KRR) as an efficient kernel-based alternative for learning high-capacity Hopfield networks. KRR utilizes the kernel trick and predicts bipolar states via regression, crucially offering a non-iterative, closed-form solution for learning dual variables. We evaluate KRR and compare its performance against Hebbian, LLR, and KLR. Our results demonstrate that KRR achieves state-of-the-art storage capacity (reaching a storage load of 1.5) and noise robustness, comparable to KLR. Crucially, KRR drastically reduces training time, being orders of magnitude faster than LLR and significantly faster than KLR, especially at higher storage loads. This establishes KRR as a potent and highly efficient method for building high-performance associative memories, providing comparable performance to KLR with substantial training speed advantages. This work provides the first empirical comparison between KRR and KLR in the context of Hopfield network learning.
@@ 1 Introduction
Hopfield networks [1] stand as a cornerstone in the study of recurrent neural networks and associative memory. They provide a simple yet powerful model for content-addressable memory, where stored patterns correspond to attractors of the network dynamics. However, the classical Hebbian learning rule [2] severely limits the storage capacity $(\beta=P/N)$ to a pattern-to-neuron ratio of $\beta\approx 0.14$, leading to the emergence of spurious states and recall failure at higher loads.
While originally proposed in the 1980s, attractor-based memory models have seen a resurgence of interest, partly due to their theoretical connections to the attention mechanisms in modern Transformers [3] and their potential for energy-efficient, brain-inspired computing hardware. This renewed interest motivates the development of learning algorithms that can overcome the classical capacity limitations and create robust, high-capacity associative memories.
To this end, research has shifted towards supervised learning paradigms. Linear methods like Linear Logistic Regression (LLR) [4] offer moderate improvements but are fundamentally constrained by the linear separability of patterns. Non-linear approaches, particularly those employing the kernel trick, provide a more powerful alternative. By implicitly mapping patterns into high-dimensional feature spaces, kernel methods can effectively handle complex, non-linear relationships.
Kernel Logistic Regression (KLR), a well-established method in statistical learning [5], has recently been demonstrated as a highly effective learning algorithm for Hopfield networks [6]. Our previous work [6] showed that KLR dramatically enhances both storage capacity and noise robustness, achieving state-of-the-art recall performance. However, KLR relies on iterative optimization, which can be computationally expensive and time-consuming, particularly for large datasets, limiting its practical scalability.
In this paper, we propose Kernel Ridge Regression (KRR) [7] as a powerful and highly efficient alternative. KRR frames the learning problem as a regression task and, crucially, offers a non-iterative, closed-form solution for its parameters. This suggests a substantial advantage in training speed. A key advantage of KRR is that the optimal parameters (dual variables) can be obtained via a direct, analytical closed-form solution, eliminating the need for computationally expensive iterative optimization.
Our primary goal is to empirically investigate whether KRR can match the state-of-the-art recall performance (capacity and noise robustness) of KLR while offering significantly faster training. We conduct a comprehensive experimental evaluation comparing KRR against Hebbian learning, LLR, and KLR. This study presents, to the best of our knowledge, the first direct empirical comparison between KRR and KLR for training Hopfield networks, highlighting the trade-offs between recall performance and learning efficiency. Our findings establish KRR as a potent and practical method for building high-performance associative memories.
@@ 2.1 Model Setup
We consider a network of $N$ bipolar neurons $\{-1,1\}^{N}$. Let $\{\boldsymbol{\xi}^{\mu}\}_{\mu=1}^{P}$ be the set of $P$ bipolar patterns to be stored in the network’s associative memory. We define a pattern matrix $\mathbf{X}\in\mathbb{R}^{P\times N}$ where the $\mu$-th row is the pattern vector $(\boldsymbol{\xi}^{\mu})^{\top}$. For supervised learning approaches requiring binary targets (LLR and KLR), we transform these patterns into target vectors $\mathbf{t}^{\mu}\in\{0,1\}^{N}$ where $t_{i}^{\mu}=(\xi_{i}^{\mu}+1)/2$. For KRR, the original bipolar values $\xi_{i}^{\mu}$ are used as targets. The network state at discrete time $t$ is represented by the vector $\mathbf{s}(t)\in\{-1,1\}^{N}$. This state evolves over subsequent time steps according to the dynamics determined by the specific learning algorithm employed (detailed in Sec. 2.2 and 2.3).
@@ 2.2.1 Baselines
  • Hebbian Learning: The standard Hebbian rule [1, 2] computes the weight matrix (diagonals zeroed)
    $\mathbf{W}^{\text{Heb}}=\frac{1}{N}\mathbf{X}^{\top}\mathbf{X}.$
    Recall uses $s_{i}(t+1)=\text{sign}(\sum_{j\neq i}\mathbf{W}_{ij}^{\text{Heb}}\mathbf{s}_{j}(t))$ This serves as the classical baseline.
  • Linear Logistic Regression (LLR): A linear supervised baseline [4], where weights $\mathbf{W}^{\text{LLR}}$ are learned iteratively via gradient descent to optimize logistic loss with a linear predictor for predicting neuron states. Recall uses the same update rule as Hebbian, with $\mathbf{W}^{\text{LLR}}$.
  • Kernel Logistic Regression (KLR): An iterative kernel-based method shown to achieve high storage capacity [6]. It learns dual variables $\boldsymbol{\alpha}^{\text{KLR}}$ by iteratively optimizing regularized logistic loss. Recall involves kernel evaluations (Sec. 2.3). We use results from [6] for comparison.
@@ 2.3 Recall Process
  • Hebbian and LLR: $s_{i}(t+1)=\text{sign}(\sum_{j\neq i}\mathbf{W}_{ij}s_{j}(t))$ (using $\mathbf{W}^{\text{Heb}}$ or $\mathbf{W}^{\text{LLR}}$), where $s_{i}(t)$ is the $i$-th element of the state $\mathbf{s}(t)$.
  • KLR and KRR: Recall operates without an explicit $N\times N$ weight matrix $\mathbf{W}$, instead utilizing the learned $P\times N$ dual variable matrix $\boldsymbol{\alpha}$ and the stored patterns $\mathbf{X}$. The update rule involves three steps, performed synchronously for all neurons:
@@ 3.1 Experimental Setup
All simulations were performed using networks composed of $N=500$ bipolar neurons, where each neuron’s state is either +1+1 or 1-1. This network size ($N=500$) was chosen to be large enough to observe the scaling behavior of different algorithms and small enough to allow for extensive simulations across varying parameters within a reasonable computational time frame on standard hardware. The patterns $\xi^{\mu}\in\{-1,1\}^{N}$ to be stored were generated randomly, with each component $\xi^{\mu}_{i}$ chosen independently and uniformly from $\{-1,1\}$. The probability $P(\xi^{\mu}_{i}=1)=0.5$. This choice of random, unbiased patterns represents a standard benchmark for evaluating the storage capacity of the Hopfield network.
We implemented and compared four distinct learning algorithms:
  • Hebbian: The classical standard Hebbian rule [1, 2], representing the traditional, non-supervised baseline. Its learning complexity is $O(N^{2})$ for computing the weight matrix.
  • LLR: A linear supervised learning method [4], serving as a stronger linear baseline that aims to make stored patterns stable fixed points. LLR learns linear classifiers for each neuron iteratively via gradient descent.
  • KLR: An advanced kernel-based supervised learning method [4], known for its high capacity and robustness, implemented based on our previous work for direct comparison. KLR learns dual variables via iterative optimization.
  • KRR: Our proposed method (detailed in Section 2.2.2), a non-iterative, closed-form kernel-based supervised learning approach.
[…]
For the iterative learning methods, LLR and KLR, a fixed learning rate $\eta=0.1$ was used. The number of learning updates was set to 100 iterations for LLR and 200 iterations for KLR based on preliminary experiments showing convergence. These numbers were chosen to ensure sufficient convergence of the respective optimization processes based on observations from our previous study [6] and preliminary runs, while also providing a practical comparison point for learning times. Using more iterations might slightly improve the performance of LLR/KLR but would further increase their learning time disadvantage relative to KRR.
Recall simulations were performed by initializing the network state $\mathbf{s}(0)$ and updating it synchronously for $T=25$ steps using the dynamics defined by each learned model (Section 2.3). This number of steps was found to be sufficient for the network state to converge to a fixed point or a limit cycle in most cases. Recall success for a stored pattern $\xi^{\mu}$ was defined as the final state $\mathbf{s}(T)$ having a normalized overlap $m(T)=\frac{1}{N}\sum_{i=1}^{N}s_{i}(T)\xi^{\mu}_{i}$ greater than $0.95$.
All simulations were implemented in Python, utilizing standard libraries such as NumPy for numerical computations. Experiments were run on a standard desktop CPU (specifically, an Intel Core i9 processor).
@@ 3.2 Storage Capacity and Noise Robustness Evaluation
Figure 1: Recall success rate vs. storage load ($\beta=P/N$) for Hebbian, LLR, KLR and KRR ($N=500$). Note: The performance curve for KLR is virtually identical to KRR.
We first assessed the storage capacity of each learning algorithm by measuring the recall success rate as a function of the storage load $\beta=P/N$. For this evaluation, recall was initiated from a clean state corresponding to each stored pattern, i.e., $\mathbf{s}(0)=\xi^{\mu}$, testing the network’s ability to maintain the learned patterns as stable fixed points. The storage load $\beta$ was varied from $0.05$ to $1.5$ in increments of $0.05$. For each value of $\beta$, $P=\lfloor\beta N\rfloor$ random patterns were generated and stored, and the recall success rate was computed as the percentage of these $P$ patterns that were successfully recalled according to our criterion ($m(T)>0.95$).
Figure 1 plots the recall success rate against the storage load $\beta$ for $N=500$ neurons. As anticipated from classical results, the standard Hebbian network exhibits a sharp phase transition and its performance drops significantly around the theoretical limit of $\beta\approx 0.14$ [2], quickly falling to 0% success rate as $\beta$ increases further. LLR significantly improves upon this, maintaining a high, near-perfect recall rate up to approximately $\beta\approx 0.85$, but it fails completely as the load approaches $\beta=0.95$ and beyond. In stark contrast, both KLR, consistent with our previous findings [6], and the proposed KRR demonstrate substantially higher storage capacity. Both kernel methods achieved a remarkable $100\%$ recall success rate across the entire tested range of storage loads, including loads exceeding $\beta=1.0$ (up to $\beta=1.5$ shown in Figure 1). This empirical result strongly indicates their capability to reliably store and retrieve patterns even in highly overloaded regimes where $P>N$, a significant improvement over linear methods.
Next, we evaluated the noise robustness of the learned memories by examining the size of the basins of attraction. We fixed the storage load at an intermediate level, $\beta=0.2$ (corresponding to $P=100$ for $N=500$ neurons), where both LLR and kernel methods typically succeed from clean initial states, but Hebbian learning already shows poor performance. For each stored pattern $\xi^{\mu}$, we generated 10 different corrupted initial states $\mathbf{s}(0)$ by flipping a certain percentage of bits relative to $\xi^{\mu}$ to achieve a desired initial overlap $m(0)$. The initial overlap $m(0)=\frac{1}{N}\sum_{i=1}^{N}s_{i}(0)\xi^{\mu}_{i}$ was varied from $0.0$ to $1.0$ in increments of $0.05$. We then measured the mean final overlap $m(T)$ achieved after $T=25$ recall steps, averaged over all $P=100$ stored patterns and the 10 trials per pattern.
Figure 2 shows the mean final overlap $m(T)$ as a function of the initial overlap $m(0)$ at $\beta=0.2$. As shown, the Hebbian network fails to converge consistently to the target pattern regardless of the initial overlap within this range, with mean final overlap remaining low. LLR shows improved robustness, successfully recalling the pattern (mean $m(T)\approx 1.0$) if the initial overlap $m(0)$ is above approximately $0.4$. In contrast, both KLR and KRR exhibit significantly enhanced robustness. They achieve perfect recall (mean $m(T)\approx 1.0$) even when starting from much noisier states, requiring an initial overlap $m(0)$ of only about $0.2$. This demonstrates that both kernel methods create considerably larger basins of attraction compared to LLR and, especially, the Hebbian rule, making them more resilient to input noise and pattern corruption. Notably, the robustness profiles of KLR and KRR appear virtually identical under these conditions, reinforcing the observation from the storage capacity test that their recall performance characteristics are highly similar.
@@
Figure 2: Noise robustness comparison: Final Overlap $m(T)$ vs. Initial Overlap $m(0)$ at $\beta=0.2$ for Hebbian, LLR, KLR and KRR. Note: The performance curve for KLR is virtually identical to KRR.
@@ 3.3 Learning Time Comparison
To quantitatively assess the practical efficiency trade-offs alongside the performance benefits, we measured and compared the learning phase computational time for LLR, KLR, and KRR across various storage loads ($\beta$). The learning time was measured as the wall-clock time using Python’s ‘time.time()’ function, capturing the total duration required to complete the learning process for each method, including any setup and optimization steps. Measurements were performed on the standard desktop CPU mentioned in Section III.A. The results, averaged over 3 independent trials for each $\beta$ value, are presented in Table 1 and visualized in Figure 3. Note that the time axis in Figure 3 is on a logarithmic scale to effectively display differences spanning multiple orders of magnitude.
Hebbian learning, being non-iterative and computationally simple ($O(N^{2})$ for weight matrix calculation), completes extremely quickly, typically in milliseconds for $N=500$. Its learning time is omitted from Table 1 and Figure 3 for clarity, as it is orders of magnitude faster than the supervised methods.
@@
The timing results reveal significant differences in efficiency and scaling among the supervised methods. LLR, requiring 100 iterative updates of gradient descent on the network weights, exhibited the slowest performance. Its learning time increased approximately linearly with the storage load $\beta$ (and thus $P$), ranging from around 3 seconds at $\beta=0.1$ to around 30 seconds at $\beta=1.0$. This scaling is consistent with the need to iterate and process information related to $P$ patterns across $N$ neurons in each update.
@@
KLR, using 200 iterative updates, was considerably faster than LLR across all tested loads, benefiting from operating in the dual space and potentially faster convergence properties for the logistic loss. Its learning time, however, still showed a clear dependence on $\beta$, increasing from around 0.07 seconds at $\beta=0.1$ to around 1.35 seconds at $\beta=1.0$. This increase is likely dominated by computations involving the $P\times P$ kernel matrix and matrix operations within the iterative optimization process, which scale with $P$.
@@
Crucially, KRR demonstrated outstanding learning efficiency. Leveraging its non-iterative, closed-form solution for the dual variables (Eq. 1), KRR consistently achieved the fastest learning times by a large margin across all tested $\beta$ values. Its learning time remained under 0.1 seconds even at the maximum load of $\beta=1.0$ (specifically, approximately 0.07 seconds as shown in Table 1), showing only a very modest increase with $\beta$. This time is primarily composed of the one-time construction of the $P\times P$ kernel matrix ($O(P^{2}N)$ or $O(P^{2})$ depending on implementation) and the subsequent solving of the $P\times P$ linear system ($O(P^{3})$ or faster).
At $\beta=1.0$, where KLR and KRR achieve peak capacity, KRR was approximately 20 times faster than KLR (with 200 updates) and over 400 times faster than LLR (with 100 updates). This dramatic speed advantage conferred by its closed-form solution is a key finding of this work, highlighting KRR’s practical superiority for rapid model training compared to iterative supervised methods, particularly in the high-capacity regime. While the recall complexity $O(PN)$ remains a shared consideration for both KLR and KRR (Section 2.3), the learning time advantage of KRR is substantial.